[figure: 9.7 cm]

ORIGIN OF MAGNETORESISTANCE ANOMALIES IN ANTIFERROMAGNETIC YBa_{2}Cu_{3}O_{6+x}

A.S. Moskvin, Yu.D. Panov Ural State University, 620083, Ekaterinburg, Russia

The crystals of tetragonal antiferromagnetic Y Ba_{2}Cu_{3}O_{6+x} (x ∼ 0.3) demonstrate an unusual behavior of the in-plane magnetoresistance, Δρ_{ab}/ρ_{ab} ∝ cos 2ϕ on the orientation of the external magnetic field within the CuO_{2} (a, b) plane [1]. This specific d-wave angular dependence is assigned to the effective hole transport through the low-lying excited purely oxygen doublet O2pe_{u} state in CuO_{4} plaquette, rather than the ground b_{1g}(d_{x^{2}−y^{2}}) state. The energy separation between these two states is estimated from the photoemission to be about 1.5 ÷ 2.0 eV, however, this is typical for lattice response accompanying the Franck-Condon transition, and does not take into account the lattice relaxation energy which magnitude is believed to be of the order of 1 ÷ 2 eV. Thus, the ”thermal” energies of additional b_{1g}(d_{x^{2}−y^{2}}) and O2pe_{u} holes forming the terms b_{1g}^{2}^{1}A_{1g} and b_{1g}e_{u} : ^{3}E_{u}, respectively, are likely to be comparable in value. The external magnetic field is believed to determine the orientation of the strong exchange field for the spin-triplet b_{1g}e_{u} : ^{3}E_{u} state of the hole CuO_{4} center and owing to the spin-orbital coupling result in a specific orbital polarization of the E_{u} doublet. Namely this gives rise to the spatial d-wave like anisotropy of the hole transport. The experimental data allow to estimate the parameter of the effective spin Hamiltonian.

[1] Y. Ando, A.N. Lavrov, K. Segawa, Phys. Rev. Lett. 83, 2813 (1999)

[figure: 13.4 cm]

Subject category :

1. Correlated Electrons and High Temperature Superconductors

Presentation mode : oral

Corresponding author : Yu.D. Panov

Address for correspondence : Lenin str. 51, Ural State University, 620083, Ekaterinburg, Russia

Email address : yuri.panov@usu.ru